OpenStudy (anonymous):

Which identity would I use to find the exact value of tan(pi/8)

OpenStudy (anonymous):

tanx=2tan(x/2)/(1-tan^2(x/2))

OpenStudy (anonymous):

where x=pi/4

OpenStudy (anonymous):

But wouldn't that be solving for x=pi/4 instead of x=pi/8 ? And what is the name of that identity, I can't find it on my identities sheet

OpenStudy (anonymous):

nah, we know the value of tanpi/4, so we will get a quadratic for tanpi/8, whose 1 root is positive nd other negative, we will reject negative by seeing the quadrant

OpenStudy (anonymous):

this is called half angle identity

OpenStudy (turingtest):

how about\[\tan(a\pm b)={\tan a\pm\tan b\over1\mp\tan a\tan b}\]??

OpenStudy (anonymous):

sir turning test, what will u put a nd b?

OpenStudy (anonymous):

\[ 1=\frac{2 u}{1-u^2} \] Solve the above for \[ \tan (\pi/4)

OpenStudy (anonymous):

Solve the above for \[ \tan (\pi/4)\]

OpenStudy (turingtest):

I'm no sir :) I was just brainstorming...

OpenStudy (turingtest):

I guess you guys are right, news to me

OpenStudy (anonymous):

elias: Would I plug in pi/4 for u?

OpenStudy (anonymous):

ok sorry sir, ( it is out of habit, i cant help calling someone sir or mam,) u nd i both will have to live with it

OpenStudy (anonymous):

In solving for u you get \[ \begin{array}{c} u=-1-\sqrt{2} \\ u=-1+\sqrt{2} \\ \end{array} \] Which one do you choose?

OpenStudy (anonymous):

If you put \( x = \frac \pi 4\) in \[ \tan (x)=\frac{2 \tan \left(\frac{x}{2}\right)}{1-\tan ^2\left(\frac{x}{2}\right)} \]

OpenStudy (anonymous):

put now \( u =\tan(\frac x 2) \) you get \[ 1=\frac{2 u}{1-u^2} \] and you finish as above.

OpenStudy (anonymous):

\[ 1-u^2=2u\\ u^2 + 2 u -1 =0 \]

OpenStudy (anonymous):

Solving it you get \[ \begin{array}{c} u=-1-\sqrt{2} \\ u=-1+\sqrt{2} \\ \end{array} \] You choose the seconf one. Why?

OpenStudy (anonymous):

\[ u=\tan(\pi/8)=-1+\sqrt{2} \]

OpenStudy (anonymous):

Ok, I'm going to work through this one, thanks for the explanation elias

OpenStudy (anonymous):

@TuringTest I am goin with him